OpenStudy (anonymous):

Write the general form of the equation of each circle. Center at the point (4,5), and containing the point (7,6).

OpenStudy (anonymous):

\[(x-4)^2+(y-5)^2=r^2\] all you need is \(r^2\)

OpenStudy (anonymous):

that is the square of the distance between the points \((4,5)\) and \((7,6)\) i.e. \[r^2=(7-4)^2+(6-5)^2\]

OpenStudy (anonymous):

Oh! Thank you!!!! :DDDDD